OpenStudy (anonymous):

the length of a rectangle is 2 meters more than its width. if the area is 63 square meters, what are the length and width?

OpenStudy (anonymous):

length=2+ width now area of rectangle= (length)(width)= 63

OpenStudy (anonymous):

\[l=2+w\] \[A=lw=63\]

OpenStudy (anonymous):

(2+width)*width=63 w^2+2w-63=0 can you do it now?

OpenStudy (anonymous):

(w+9)(w-7)=0 w=9,7

OpenStudy (anonymous):

@RenaeMonroe ??

OpenStudy (anonymous):

9*7 =63 and 7+2=9

OpenStudy (anonymous):

how did you get 9,7 ???

OpenStudy (anonymous):

by solving and factorizing \[w^2+2w-63=0\]

OpenStudy (anonymous):

\[w^2+9w-7w-63=(w+9)(w-7)\]

OpenStudy (anonymous):

k i see

OpenStudy (anonymous):

hope that helps you

OpenStudy (anonymous):

it has thank you let me see if i can find the length and width then i will post it and you can tell me if it is right or wrong

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

ok im lost i cant figure out what to do

OpenStudy (anonymous):

see you got w=-9,7 so -9 can't be a length therefore 7=width

OpenStudy (anonymous):

where did you get -9

OpenStudy (anonymous):

see we have (w+9)(w-7)=0 therefore w =-9 and 7

OpenStudy (anonymous):

ok so i have a=l*w 63=(w+2)*w 63=w^2+2w 0=w^2+2w-63 0=(w-7)(w+9) w-7=0 or w+9=0 w=7 or w=-9 but -9 wont work because you can not have a negative length so the only right answer is w=7 is this right

OpenStudy (anonymous):

yes correct

OpenStudy (anonymous):

thank you for helping

OpenStudy (anonymous):

:) YW